OpenStudy (anonymous):

List the critical numbers of the following function in increasing order. f(x)=3x^3+3x^2−3x

OpenStudy (anonymous):

Take the derivative, find where the derivative is equal to 0 or does not exist.

OpenStudy (anonymous):

so 9x^2+ 6x-3. how do i find where it is = 0 or DNE?

OpenStudy (accessdenied):

Set your function equal to zero and solve. Are you familiar with solving quadratic equations? \(9x^2 + 6x - 3 = 0\) Solve for the x-values.

OpenStudy (accessdenied):

There are a few strategies for solving quadratics. I will describe my method: a) Try to factor the equation. If it is factorable, then we can use the property that A*B = 0 means A = 0 or B = 0 and solve two linear equations. b) If it is not factorable or unclear if it even factors, we can use the quadratic formula. For an equation in the form \(ax^2 + bx + c = 0\), we have: \[ \qquad x = \frac{\neg b \pm \sqrt{b^2 - 4ac}}{2a} \]As the solutions.